OpenStudy (anonymous):

What is the new volume of a rectangular prism that is originally 18 cubic meters and is then resized using a scale factor of 2 on each side length?

OpenStudy (anonymous):

let a b and c be the three sides volume, a*b*c =18 when resized sides are multiplied by 2 2a *2b *2c= 8*abc =8 *18

OpenStudy (anonymous):

So 8*18 will be my answer?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Thank you :)